[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jul 09, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Jul 9 18:01:11 UTC 2016


HSU Shift Log for Sat Jul 09 2016 14:01:05 GMT-0400 (EDT)

Shift: H

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues: For Celia, SAB produced the regular synoptic hour fix at 1200  
Z and then sent a corrected version of the 12Z fix (using a microwave pass)  
a couple of hours later at 14 Z. In the ATCF, it appears that the corrected  
fix overwrote the one done in real time, which was not a true real-time  
assessment.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160709/4e4c61e6/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list